# Texture-aware and Shape-guided Transformer for Sequential DeepFake Detection

### v2

@@ Abstract
Sequential DeepFake detection is an emerging task that aims to predict the manipulation sequence in order. Existing methods typically formulate it as an image-to-sequence problem, employing conventional Transformer architectures for detection. However, these methods lack dedicated design and consequently result in limited performance. In this paper, we propose a novel Texture-aware and Shape-guided Transformer to enhance detection performance. Our method features four major improvements. Firstly, we describe a texture-aware branch that effectively captures subtle manipulation traces with the Diversiform Pixel Difference Attention module. Then we introduce a Bidirectional Interaction Cross-attention module that seeks deep correlations among spatial and sequential features, enabling effective modeling of complex manipulation traces. To further enhance the cross-attention, we describe a Shape-guided Gaussian mapping strategy, providing initial priors of the manipulation shape. Finally, observing that the latter manipulation in a sequence may influence traces left in the earlier one, we intriguingly invert the prediction order from forward to backward, leading to notable gains as expected. Extensive experimental results demonstrate that our method outperforms others by a large margin, highlighting the superiority of our method.
@@ 1 Introduction
Recent advancements in deep generative techniques have significantly improved the visual quality of generated faces. One such technique, DeepFake, allows for effortless manipulation of faces, with results that are often imperceptible to the naked eye. This technique has found applications in virtual reality and movie special effects, but it also raises severe security concerns He et al. (2021). To address these concerns, various forensic methods have been proposed to expose DeepFakes, garnering considerable attention.
@@
Conventional forensics methods mainly focus on determining whether a given face image is authentic or forged, which typically falls into the binary classification task. However, with the rapid evolution of face editing tools, the manipulation process has become more versatile and user-friendly, enabling the use of multiple specific operations iteratively until the desired effects are reached. Under this circumstance, existing methods have limitations as they can only identify the final step of manipulation, providing limited insights into the manipulation details and hindering traceability and interpretability. To overcome this issue, a new forensics task called Sequential DeepFake Detection Shao et al. (2022) has emerged recently, aiming to uncover the sequences of face manipulations (see Fig. 1(a,b)). This task expands the scope of DeepFake detection and poses a new challenge to this field.
Figure 1: (a) and (b) describe the difference between conventional DeepFake detection and sequential DeepFake detection. (c) illustrates the process of classic sequential DeepFake detection methods, while (d) represents our proposed method.
Typically, Transformer-based architectures are employed in sequential DeepFake detection to capture traces of different manipulations and predict the manipulation sequences, following the pipeline of image-to-sequence Vinyals et al. (2015). These methods commonly employ a plain Transformer encoder to extract spatial clues and utilize a Transformer decoder to capture the sequential clues (see Fig. 1(c)). Despite these methods showing a certain effectiveness, many critical perspectives are lacking in further exploration.
  1. $\clubsuit$
    Fine-grained Traces Extraction: Generative models often exhibit imperfect semantic disentanglement. One specific manipulation may affect the entire facial structure and leave subtle traces on unintended manipulated regions. To expose the sequential manipulation, existing methods usually utilize a general Transformer encoder to capture those traces. However, this encoder concentrates on the global relationship of the entire faces and may fail to capture the subtle manipulation traces.
  2. $\clubsuit$
    Manipulation Shape Guidance: In the training stage, existing methods enhance the modeling of sequential relation in the decoder by estimating the scale and position of manipulation regions as priors. However, such priors are impractical as the manipulation regions are likely irregular, which could not be well described by only using the scale and position.
  3. $\clubsuit$
    Prediction Order: Existing methods predict the manipulation sequences in a forward order, which is the same as the actual manipulation order. However, the previous manipulation always hides beneath the subsequent manipulation. Think that it is challenging to solely take out the center of an “onion” without peeling off its skin layers. Similarly, the first manipulation is the most challenging to detect, and this challenge hinders the model from gradually capturing sequential traces.
@@ 3. $\clubsuit$
In this paper, we delve into the above perspectives and propose a Texture-aware and Shape-guided Transformer (TS-Former) to expose sequential DeepFakes (see Fig. 1(d)). Our method introduces four major improvements, each derived from a specific perspective. 1) In the encoder, we develop two branches for capturing different types of manipulation traces. One is the vanilla branch, which is composed of conventional Transformer blocks that capture global relationships. Inspired by the effect of textures in forgery detection Liu et al. (2020); Yang et al. (2021), we design a texture-aware branch to capture the subtle manipulation traces. In this branch, we describe a Diversiform Pixel Difference Attention (DPDA) module, which adaptively integrates various texture extraction operations into the self-attention mechanism, with guidance from the vanilla branch. 2) Then we propose a new Bidirectional Interaction Cross-attention (BICA) module to effectively model correlations among the spatial features from the encoder and sequential manipulation annotations. Unlike standard cross-attention, our module simultaneously processes multiple features and employs a multi-level fusion strategy for more effective correlations. 3) To enhance cross-attention, we propose a Shape-guided Gaussian Mapping strategy to provide the priors of manipulation shapes. Instead of estimating the scale and position, we predict the Gaussian distribution of their manipulated shapes using a Variational Auto-encoder (VAE). 4) Subsequently, we invert the order of the manipulation sequences in the training phase. This simple “trick” notably improves the detection performance as expected. Our method is evaluated on a public sequential DeepFake dataset and compared with several state-of-the-art methods, demonstrating the effectiveness of our method in detecting sequential DeepFakes.
The contribution of this paper is summarized as
  1. 1.
    We introduce a texture-aware branch to capture subtle manipulation traces. To achieve this, we describe a Diversiform Pixel Difference Attention (DPDA) module, which adaptively integrates various texture extraction operations into the self-attention mechanism.
  2. 2.
    We describe a Bidirectional Interaction Cross-attention (BICA) module to effectively model the correlations among spatial and sequential features in a multi-level fusion strategy. Then we propose a Shape-guided Gaussian Mapping strategy to enhance cross-attention.
  3. 3.
    By intriguingly inverting the prediction order of manipulation sequences, we achieve notable performance gains.
  4. 4.
    Through extensive experiments, we demonstrate the efficacy of our method and thoroughly study the effect of each component of our method for further exploration.
@@ 2 Related Works
DeepFake Detection. To combat the threats posed by DeepFake, a considerable amount of methods have been proposed for DeepFake detection Li et al. (2018); Yang et al. (2019); Gu et al. (2022); Xu et al. (2023); Dong et al. (2023); Guo et al. (2023); Wang et al. (2023). These methods can typically be classified into different categories in terms of the way of feature capturing. The methods of Li et al. (2018); Yang et al. (2019); Agarwal et al. (2019); Qi et al. (2020); Yu et al. (2024) expose DeepFake by capturing abnormal biological signals, such as eye blinking, behavior patterns, or heart rhythm. Other methods focus on extracting the spatial artifacts introduced either by face generation or the blending operations Li and Lyu (2019); Li et al. (2020); Zhao et al. (2021c); Shiohara and Yamasaki (2022). Instead of capturing the spatial artifacts, the methods of Li et al. (2021); Liu et al. (2021); Dzanic et al. (2020) demonstrate the effectiveness of frequency information and capture representative traces in the frequency domain. Moreover, many recent methods design dedicated models to capture the traces automatically Zhao et al. (2021a); Guo et al. (2023); Wang et al. (2023) However, these methods can only predict authenticity, which can not handle and analyze the multi-step alterations in practice.
Sequential DeepFake Detection. Sequential DeepFake detection is proposed to predict the sequential manipulations in order. This task is generally formulated as the image-to-sequence problem and utilizes sequential-related architectures for detection. The method of Shao et al. (2022) solves this task using a plain Transformer architecture. In this paper, we further explore this task and describe a new method dedicated to sequential DeepFake detection. Our method features four major improvements and has demonstrated effectiveness in experiments.
@@
3 Method
Our method consists of a Transformer encoder and a Transformer decoder. The encoder takes the face images as input and extracts spatial manipulation features, while the decoder models the sequential relation based on these features and sequence annotations to predict the facial manipulation sequence.
The encoder is composed of a vanilla branch and a new texture-aware branch. In the texture-aware branch, we introduce the Diversiform Pixel Difference Attention (DPDA) module to capture subtle manipulation traces (Sec. 3.1). For the decoder, we propose a Bidirectional Interaction Cross-attention strategy to fuse the extracted spatial features with the sequence annotation embeddings, enabling effective modeling of sequential relations (Sec. 3.2). To provide more guidance, we introduce Shape-guided Gaussian Mapping to enhance the effect of cross-attention in sequential modeling (Sec. 3.2). We then introduce the procedure for inverting the prediction orders (Sec. 3.2) and outline the overall loss functions (Sec. 3.3).
3.1 Spatial Encoder
Tokenization. Denote a face image as $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$. To tokenize this face image, we employ a simple CNN stem following He et al. (2015). Let a CNN stem be $\mathcal{F}$. The input tokens can be obtained as $\mathbf{x}^{0}=\mathcal{F}(\mathbf{I})+\mathbf{p},\mathbf{z}^{0}\in\mathbb{R}^{h\times w\times c}$, where $\mathcal{F}(\mathbf{I})$ is the feature maps from CNN model and $\mathbf{p}$ is a positional embedding. Then we flatten the spatial dimension of these tokens to one dimension as the input of the encoder.
Vanilla Branch. This branch contains several Transformer blocks and each block calculates a multi-head self-attention. Denote the index of a block as $l$ and the input token of this block as $\mathbf{x}^{l}$. Thus the operation of this block can be written as
$\hat{\mathbf{x}}^{l}$ $=\mathbf{x}^{l}+{\rm MSA}({\rm LN}(\mathbf{x}^{l})),$ (1)
$\hat{\mathbf{x}}^{l+1}$ $=\hat{\mathbf{x}}^{l}+{\rm FFN}({\rm LN}(\hat{\mathbf{x}}^{l})),$
where ${\rm MSA}$ denotes the multi-head self-attention operation, ${\rm LN}$ denotes the LayerNorm Ba et al. (2016), FFN is the feed-forward network, which consists of two FC layers with GELU Hendrycks and Gimpel (2016).
Let $\mathbf{x}$ be the input tokens of MSA layer. In each head, three projection matrices are applied to generate query, key, and value features. Then the self-attention is conducted over the input tokens spatially to learn the coarse-grained traces. Within the $i$-th head, the self-attention can be defined as
$Q_{i}=\mathbf{x}W^{Q}_{i},K_{i}=\mathbf{x}W^{K}_{i},V_{i}=\mathbf{x}W^{V}_{i},$ (2)
$h_{i}={\rm Softmax}(Q_{i}K^{\top}_{i}/\sqrt{d_{i}})V_{i},$
@@ 3.1 Spatial Encoder
where $W^{Q}_{i},W^{K}_{i},W^{V}_{i}$ are the projection matrices, $d_{i}$ is the number of hidden dimensions, and $h_{i}$ is the self-attention of $i$-th head. Then we concatenate all these heads and perform a linear transformation with $W^{o}$ to obtain the final output.
Texture-aware Branch. To extract these subtle manipulation traces, we develop a new Transformer block designed specifically for extracting texture features. Our block draws inspiration from central difference operators Yu et al. (2020b); Juefei-Xu et al. (2017), which are convolutions used for detecting local textures. In our method, we derive and extend their essence from the convolution style to the self-attention mechanism, and propose a Diversiform Pixel Difference Attention (DPDA) module that captures fine-grained spatial traces while retaining the global view within the vanilla self-attention.
1) Revisit of Central Difference Convolution. Central difference convolution is a special case of convolution, which emphasizes the center-oriented gradients in the local receptive field region. Given a feature map $\mathbf{x}$, this operation can be formulated as
$\mathbf{x}^{\prime}(p_{0})=$ $\sum_{p_{n}\in\mathcal{O}}\mathbf{w}(p_{n})\cdot(\mathbf{x}(p_{0}+p_{n})-\mathbf{x}(p_{0})),$ (3)
2) Diversiform Pixel Difference Attention. Inspired by the central difference operations, we extend these operations and integrate such spirit into the self-attention mechanism following Yu et al. (2020b, a); Miao et al. (2023). Specifically, given the feature map $\mathbf{x}$, we first perform a vanilla convolution on it to generate the query feature, which is defined as
$\mathbf{x}^{Q}(p_{0})=$ $\sum_{p_{n}\in\mathcal{O}}\mathbf{w}(p_{n})\cdot\mathbf{x}(p_{0}+p_{n}).$ (4)
Then we would like to capture the fine-grained spatial features as the source of key and value features. Since the central difference operations focus on the center-oriented gradients, they may overlook other texture patterns, which hinders the effectiveness of capturing manipulation traces. To obtain a diverse range of gradient information, we leverage the concept of Extended Local Binary Pattern (ELBP), which can encode the pixel relations from various perspectives, including central, angular, and radial directions Su et al. (2021). The angular direction calculates the difference between adjacent pixels in the outline of the kernel in a certain order, while the radial direction calculates the difference between radial adjacent pixels, see Fig. 2. To obtain the key and value features, we adaptively ensemble these operations using a learnable coefficient vector, which is formulated as
$\mathbf{x}^{K,V}(p_{0})=\sum_{p_{n}\in\mathcal{O}}\mathbf{w}(p_{n})\cdot$ $\left(\alpha_{c}\cdot\underbrace{(\mathbf{x}(p_{0}+p_{n})-\mathbf{x}(p_{0}))}_{\mathcal{D}_{c}}\right.$ (5)
$+\alpha_{a}\cdot\underbrace{(\mathbf{x}(p_{0}+p_{n})-\mathbf{x}(p_{n-1}))}_{\mathcal{D}_{a}}$
$+\alpha_{r}\cdot\left.\underbrace{(\mathbf{x}(p_{0}+p_{n})-\mathbf{x}(p_{\phi(n)}))}_{\mathcal{D}_{r}}\right)$
$=\sum_{p_{n}\in\mathcal{O}}\mathbf{w}(p_{n})\cdot$ $\left(\underbrace{\left[\alpha_{c},\alpha_{a},\alpha_{r}\right]}_{\bm{\alpha}}\left[\begin{array}[]{l}\mathcal{D}_{c}\\ \mathcal{D}_{a}\\ \mathcal{D}_{r}\end{array}\right]\right).$
In this equation, $\mathcal{D}_{c},\mathcal{D}_{a},\mathcal{D}_{r}$ denote central, angular, and radial difference operations respectively. We use $3\times 3$ as the local receptive field $\mathcal{O}$ for instance and set $\mathcal{O}=\{(-1,-1),(-1,0),...,(1,1)\}$. Thus $\phi(n)$ in radial difference operation can defined as $\phi(n)=(2x,2y)$ if $n=(x,y)$. $\bm{\alpha}$ is the coefficient vector to balance the contribution of each operation.
Given that different DeepFake face images may contain different manipulation traces, we design a learning scheme to adaptively balance the contribution strength of different DPDA operations. As our first branch concentrates on extracting coarse-grained information, its intermediate features can provide global instruction on the composition of manipulations. Denote the feature map of the corresponding block in our coarse-grained branch as $\mathbf{x}^{\prime}$. To achieve this adaptive learning, we design a sub-network $\mathcal{H}$ that contains several convolution layers. This network learns the $\bm{\alpha}$ based on $\mathbf{x}^{\prime}$ as
$\bm{\alpha}={\rm Softmax}(\mathcal{H}(\mathbf{x}^{\prime})).$ (6)
After obtaining $\mathbf{x}^{Q}$ and $\mathbf{x}^{K,V}$, we project the query feature $\mathbf{x}^{Q}$ into query tokens, and project $\mathbf{x}^{K,V}$ into key and value tokens. Within each head, ${Q}_{i}=\mathbf{x}^{Q}{W}^{Q}_{i},{K}_{i}=\mathbf{x}^{K,V}{W}^{K}_{i},{V}_{i}=\mathbf{x}^{K,V}{W}^{V}_{i}$. Then the multi-head self-attention is calculated using the Eq. (2).
3) Diversiform Pixel Difference Transformer Block. This block integrates the Diversiform Pixel Difference Attention (DPDA) module, which is similar to the Eq. (1). The overall process can be written as
$\hat{\mathbf{x}}^{l}$ $=\mathbf{x}^{l}+{\rm DPDA}({\rm LN}(\mathbf{x}^{l})),$ (7)
$\hat{\mathbf{x}}^{l+1}$ $=\hat{\mathbf{x}}^{l}+{\rm FFN}({\rm LN}(\hat{\mathbf{x}}^{l})),$
With this strategy, this branch can effectively extract subtle manipulation traces, enhancing the performance in sequential detection.
3.2 Sequential Decoder
In order to predict sequential manipulations, we develop a sequential decoder architecture, which takes as input the features from the encoder and the embeddings of manipulation sequences. To effectively integrate these types of features into the decoder, we propose a Bidirectional Interactive Cross-attention (BICA) strategy that can properly integrate the features from the encoder into the decoder. Moreover, to improve the capacity of cross-attention, we introduce a shape-guided Gaussian mapping strategy to enhance its effectiveness.
Spatial and Sequential Feature Fusing. To fully leverage the benefits of both coarse-grained and fine-grained features, we propose a Bidirectional Interaction Cross-attention (BICA) module. This module employs a hierarchical attention mechanism to explore the correlation between features and sequential annotations.
1) Revisit of Conventional Cross-attention The cross-attention is typically employed in the decoder of the transformer for sequence tasks. Unlike self-attention mechanisms, the query, and key and value features usually originate from different sources. Given two input tokens from different sources $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$, the multi-head cross-attention (MCA) can be defined as

### v4

# Texture, Shape and Order Matter: A New Transformer Design for Sequential DeepFake Detection

@@ Abstract
Sequential DeepFake detection is an emerging task that predicts the manipulation sequence in order. Existing methods typically formulate it as an image-to-sequence problem, employing conventional Transformer architectures. However, these methods lack dedicated design and consequently result in limited performance. As such, this paper describes a new Transformer design, called TSOM, by exploring three perspectives: Texture, Shape, and Order of Manipulations. Our method features four major improvements: ❶ we describe a new texture-aware branch that effectively captures subtle manipulation traces with a Diversiform Pixel Difference Attention module. ❷ Then we introduce a Multi-source Cross-attention module to seek deep correlations among spatial and sequential features, enabling effective modeling of complex manipulation traces. ❸ To further enhance the cross-attention, we describe a Shape-guided Gaussian mapping strategy, providing initial priors of the manipulation shape. ❹ Finally, observing that the subsequent manipulation in a sequence may influence traces left in the preceding one, we intriguingly invert the prediction order from forward to backward, leading to notable gains as expected. Extensive experimental results demonstrate that our method outperforms others by a large margin, highlighting the superiority of our method.
@@ 1 Introduction
Recent advancements in deep generative techniques have significantly improved the visual quality of generated faces. One such technique, DeepFake, allows for effortless manipulation of faces, with results that are often imperceptible to the naked eye. This technique has found applications in virtual reality and movie special effects, but it also raises severe security concerns [9]. To address these concerns, various forensic methods have been proposed to expose DeepFakes, garnering considerable attention [15, 34, 14].
@@
Conventional forensics methods mainly focus on determining whether a given face image is authentic or forged, which typically falls into the binary classification task (see Fig. 1(a)). However, with the rapid evolution of face editing tools, the manipulation process has become more versatile and user-friendly, enabling the use of multiple specific operations iteratively until the desired effects are reached. In this context, existing methods can only identify the final step of manipulation, providing limited insights into the manipulation details. To overcome this limitation, a new forensics task called Sequential DeepFake Detection (SDD) [23] has emerged recently. In contrast to the conventional DeepFake detection (i.e., One-step DeepFake Detection (ODD)), SDD attempts to uncover the sequences of face manipulations in order. For example, users might create a DeepFake face by sequentially editing regions like hair, eyes, and lips. Given such a DeepFake face, SDD can predict the manipulation sequence as “Hair-Eye-Lip” (see Fig. 1(b)). This emerging task improves the practicality of DeepFake detection in real-world scenarios, offering more comprehensive evidence for traceability and interoperability (e.g., pinpointing where, when, and what manipulations are applied). This endeavor expands the horizon of DeepFake detection and poses a fresh challenge to the field.
Figure 1: (a) and (b) correspond to One-step DeepFake Detection (ODD) and sequential DeepFake Detection (SDD). (c) illustrates the process of typical Sequential DeepFake Detection (SDD) architecture. (d) represents our proposed method TSOM. In contrast to (c), our method focuses on exploring three perspectives: Texture, Shape, and Order of Manipulations, and introduces four major improvements: a Diversiform Pixel Difference Attention (DPDA) module, a Multi-source Cross-attention (MSCA) module, a Shape-guided Gaussian Mapping (SGM) strategy and an Inverted Order Prediction (IOP).
In general, the task of SDD can be conceptualized as an image-to-sequence problem [27], where the typical solution involves leveraging Transformer-based architectures to predict the manipulation sequences [31, 23]. Concretely, a spatial encoder with conventional self-attention mechanisms is utilized to extract spatial manipulation traces. Then, a sequential decoder is employed to establish conventional cross-attention [26] with the encoder features and manipulation annotation sequence, facilitating the capture of sequential manipulation traces (see Fig. 1(c)).
Despite these methods showing certain effectiveness, many critical perspectives are lacking in further exploration: 1) Neglecting texture details: Generative models often exhibit imperfect semantic disentanglement, wherein one specific manipulation may affect the entire facial structure, leaving subtle traces on unintended manipulated regions. These subtle traces often manifest more prominently in the textures [17, 33]. However, the existing methods rely on conventional self-attention mechanisms that concentrate more on global correlations, thus overlooking fine-grained features like textures. 2) Lack of manipulation shape guidance: The existing pipeline accomplishes detection from a holistic perspective, directly predicting the manipulations without considering their specific shapes. Note that the manipulation regions often exhibit various shapes and irregularities. Incorporating knowledge about manipulation shapes can guide the network to gain a deeper understanding of the manipulations. While existing method [23] incorporates the scale and position of manipulation regions, they could not effectively characterize irregular manipulation regions. 3) Improper prediction order: Existing methods predict the manipulation sequences in a forward order, the same as the actual manipulation order. However, the subsequent manipulation usually affects the appearance of the preceding ones (see Fig. 2)), i.e., the preceding manipulation “hides beneath” the subsequent one, akin to the layers of an “onion”. The first manipulation step (e.g., Hair as shown in Fig. 1(b)) acts as the core of the onion. Extracting this core without peeling away its outer layers (e.g., Eye and lip as shown in Fig. 1(b)) would be improper.
Figure 2: The impact of the sequential manipulation on facial attributes. The second manipulation impacts the first manipulation: (left) Eye and (right) Eyeglasses.
@@
Delving into these perspectives, we propose a new Transformer called TSOM (Texture, Shape and Order of Manipulations) to expose sequential DeepFakes (see Fig. 1(d)). Our method introduces four major improvements: ❶ In the encoder, we develop two branches for capturing different types of manipulation traces. One branch, termed the vanilla branch, employs conventional Transformer blocks to capture global relationships. The other branch, a new texture-aware branch, is designed to capture subtle manipulation traces. Within this branch, we describe a Diversiform Pixel Difference Attention (DPDA) module, which dynamically integrates various texture extraction operations into the self-attention mechanism, instructed by the knowledge from the vanilla branch. ❷ Then we propose a new Multi-source Cross-attention (MSCA) module to effectively model correlations between spatial features from the encoder and sequential manipulation annotations. Unlike traditional cross-attention, our module processes multiple features simultaneously and employs a multi-level fusion strategy for more effective correlations. ❸ To further enhance cross-attention, we propose a Shape-guided Gaussian Mapping (SGM) strategy to provide priors on manipulation shapes. These priors are estimated using a Variational Auto-encoder (VAE) [11]. ❹ During training, we invert the order of the manipulation sequences. This simple “trick” notably improves the detection performance, as expected.
Our method is evaluated on a sequential DeepFake dataset [23] and compared with several state-of-the-art methods. The results demonstrate its superiority in detecting sequential DeepFakes. Furthermore, we conduct a comprehensive analysis of each component within our method, corroborating the efficacy of the proposed strategies.
@@ 2 Related Works
One-step DeepFake Detection (ODD). To combat the threats posed by DeepFake, a considerable amount of methods have been proposed for DeepFake detection [14, 34, 6, 32, 3, 7, 30]. These methods can typically be classified into different categories in terms of the way of feature capturing. The methods of [14, 34, 1, 20, 35] expose DeepFake by capturing abnormal biological signals, such as eye blinking, behavior patterns, or heart rhythm. Other methods focus on extracting the spatial artifacts introduced either by face generation or the blending operations [15, 13, 40, 24]. Instead of capturing the spatial artifacts, the methods of [12, 16, 4] demonstrate the effectiveness of frequency information and capture representative traces in the frequency domain. Moreover, many recent methods design dedicated models to capture the traces automatically [38, 7, 30]. However, these methods can only predict authenticity, which can not handle and analyze the multi-step alterations in practice.
Sequential DeepFake Detection (SDD). Sequential DeepFake detection is an emerging task focused on predicting the sequential manipulations in order. This task is typically formulated as the image-to-sequence problem, utilizing architectures tailored for sequential tasks in detection. The work of [23] is pioneering in addressing this problem, employing the Transformer architecture [26], which includes a conventional encoder and decoder with vanilla self-attention mechanisms. Subsequently, MMNet [31] improved its performance through the incorporation of multi-collaboration and multi-supervision modules. Despite these promising results, this task remains largely unexplored.
3 Method
Our method is developed on a Transformer-based architecture comprising a spatial encoder and a sequential decoder. The encoder is designed to extract spatial manipulation features, including a vanilla branch and a texture-aware branch. The vanilla branch employs several conventional Transformer blocks to capture the global correlations. In the texture-aware branch, we introduce the Diversiform Pixel Difference Attention (DPDA) module to capture subtle manipulation traces (Sec. 3.1). The decoder models the sequential relation based on the spatial features and corresponding sequence annotations to predict the facial manipulation sequence. In this decoder, we propose a Multi-source Cross-attention (MSCA) strategy to fuse the extracted spatial features with the sequence annotation embeddings, enabling effective modeling of sequential relations (Sec. 3.2). Moreover, we introduce Shape-guided Gaussian Mapping (SGM) to enhance the effect of cross-attention in sequential modeling (Sec. 3.3). We then introduce the procedure for Inverted Order Prediction (IOP) (Sec. 3.4) and outline the overall loss functions (Sec. 3.5).
3.1 Texture-aware Branch
[…]
To extract these subtle manipulation traces, we develop a new Transformer Block for extracting texture features. Our block draws inspiration from central difference operators [37, 10], which are convolutions used for detecting local textures. In our method, we derive and extend their essence from the convolution style to the self-attention mechanism, and propose a Diversiform Pixel Difference Attention (DPDA) module that captures fine-grained spatial traces while retaining the global view within the vanilla self-attention.
@@ 3.1 Texture-aware Branch
Figure 4: Overview of the sequential decoder (left). Illustration of Multi-source Cross-attention(MSCA). MCA denotes Multi-head Cross-attention.
Revisit of Central Difference Convolution. Central difference convolution is a special case of convolution, which emphasizes the center-oriented gradients in the local receptive field region. Given a feature map $\mathbf{x}$, this operation can be formulated as
$\mathbf{x}^{\prime}(p_{0})=$ $\sum_{p_{n}\in\mathcal{O}}\mathbf{w}(p_{n})(\mathbf{x}(p_{0}+p_{n})-\mathbf{x}(p_{0}))$ (1)
Diversiform Pixel Difference Attention (DPDA). Inspired by the central difference operations, we extend these operations and integrate such spirit into the self-attention mechanism following [37, 36, 19]. Specifically, given the feature map $\mathbf{x}$, we first perform a vanilla convolution on it to generate the query feature, which is defined as
$\mathbf{x}^{Q}(p_{0})=$ $\sum_{p_{n}\in\mathcal{O}}\mathbf{w}(p_{n})\mathbf{x}(p_{0}+p_{n})$ (2)
Then we would like to capture the fine-grained spatial features as the source of key and value features. Since the central difference operations focus on the center-oriented gradients, they may overlook other texture patterns, which hinders the effectiveness of capturing manipulation traces. To obtain a diverse range of gradient information, we leverage the concept of Extended Local Binary Pattern (ELBP), which can encode the pixel relations from various perspectives, including central, angular, and radial directions [25]. The angular direction calculates the difference between adjacent pixels in the outline of the kernel in a certain order, while the radial direction calculates the difference between radial adjacent pixels, see Fig. 3. To obtain the key and value features, we adaptively ensemble these operations using a learnable coefficient vector, which is formulated as
$\mathbf{x}^{K,V}(p_{0})=\sum_{p_{n}\in\mathcal{O}}\mathbf{w}(p_{n})\left(\underbrace{\left[\alpha_{c},\alpha_{a},\alpha_{r}\right]}_{\bm{\alpha}}\left[\begin{array}[]{l}\mathcal{D}_{c}\\ \mathcal{D}_{a}\\ \mathcal{D}_{r}\end{array}\right]\right)$ (3)
In this equation, $\mathcal{D}_{c},\mathcal{D}_{a},\mathcal{D}_{r}$ denote central, angular, and radial difference operations respectively, which can be defined as
$\mathcal{D}_{c}$ $=\mathbf{x}(p_{0}+p_{n})-\mathbf{x}(p_{0}),$ (4)
$\mathcal{D}_{a}$ $=\mathbf{x}(p_{0}+p_{n})-\mathbf{x}(p_{n-1}),$
$\mathcal{D}_{r}$ $=\mathbf{x}(p_{0}+p_{n})-\mathbf{x}(p_{\phi(n)})$
We use $3\times 3$ as the local receptive field $\mathcal{O}$ for instance and set $\mathcal{O}=\{(-1,-1),(-1,0),...,(1,1)\}$. Thus $\phi(n)$ in radial difference operation can defined as $\phi(n)=(2x,2y)$ if $n=(x,y)$. $\bm{\alpha}$ is the coefficient vector to balance the contribution of each operation.
Adaptive Balance. Since various DeepFake faces exhibit diverse manipulation traces, the coefficient vector $\bm{\alpha}$ should be different. Thus, we design a strategy to dynamically regulate the influence of different DPDA operations. As the vanilla branch concentrates on global correlations, it has deeper insight into the manipulation composition. Hence, we set $\bm{\alpha}$ trainable and adapt it based on the knowledge from the vanilla branch. Let $\mathbf{x}^{\prime}$ be the intermediate feature of a block within the vanilla branch. We develop a sub-network $\mathcal{H}$ to learn $\bm{\alpha}$ as
$\small\bm{\alpha}={\rm Softmax}(\mathcal{H}(\mathbf{x}^{\prime}))$ (5)
After obtaining $\mathbf{x}^{Q}$ and $\mathbf{x}^{K,V}$, we project the query feature $\mathbf{x}^{Q}$ into query tokens, and project $\mathbf{x}^{K,V}$ into key and value tokens. Within each head, ${Q}_{i}=\mathbf{x}^{Q}{W}^{Q}_{i},{K}_{i}=\mathbf{x}^{K,V}{W}^{K}_{i},{V}_{i}=\mathbf{x}^{K,V}{W}^{V}_{i}$. Then the multi-head self-attention is calculated using
$h_{i}={\rm Softmax}(Q_{i}K^{\top}_{i}/\sqrt{d_{i}})V_{i},$ (6)
@@
where $h_{i}$ is the self-attention of $i$-th head. Then we concatenate all these heads and perform a linear transformation with $W^{o}$ to obtain the final output.
3.2 Multi-source Cross-attention (MSCA)
Revisit of Conventional Cross-attention. Typically, the cross-attention [26] is employed for modeling such correlation [23]. However, these mechanisms usually process the features from only two sources $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ (e.g., the encoder features and sequential embedding). The conventional Multi-head Cross-attention (MCA) can be defined as